OpenStudy (anonymous):

Can you find the product of (0.5n^5)^2(10n^7)^3, and explain how you got the answer please?

OpenStudy (amistre64):

id assume youd need to have like bases

OpenStudy (amistre64):

\[.5^2 (n^5)^2 10^3 (n^7)^3\]

OpenStudy (amistre64):

.5^2 = .5*.5 = .25 10^3 = 10.10.10 = 1000 (n^5)^2 = n^5.n^5 = n^10 (n^7)^3 = n^7.n^7.n^7 = n^21

OpenStudy (hoblos):

=0.25n^10 * 1000n^21 =250n^31

OpenStudy (amistre64):

how to explain it tho, what part confuses you?

OpenStudy (anonymous):

Which one of you has the right answer?

OpenStudy (amistre64):

thats not an explanation seeing that we both have the same answer ....

OpenStudy (hoblos):

amistre64 didn't answer he just gave you the way so that you can solve it by tour own

OpenStudy (anonymous):

Oh I see. Thank you for your help.